# Dynamic Post-Hoc Neural Ensemblers

### v1

@@ Abstract
Ensemble methods are known for enhancing the accuracy and robustness of machine learning models by combining multiple base learners. However, standard approaches like greedy or random ensembles often fall short, as they assume a constant weight across samples for the ensemble members. This can limit expressiveness and hinder performance when aggregating the ensemble predictions. In this study, we explore employing neural networks as ensemble methods, emphasizing the significance of dynamic ensembling to leverage diverse model predictions adaptively. Motivated by the risk of learning low-diversity ensembles, we propose regularizing the model by randomly dropping base model predictions during the training. We demonstrate this approach lower bounds the diversity within the ensemble, reducing overfitting and improving generalization capabilities. Our experiments showcase that the dynamic neural ensemblers yield competitive results compared to strong baselines in computer vision, natural language processing, and tabular data.
@@ 1 Introduction
Ensembling machine learning models is a well-established practice among practitioners and researchers, primarily due to its enhanced predictive performance over single-model predictions. Ensembles are favored for their superior accuracy and ability to provide calibrated uncertainty estimates and increased robustness against covariate shifts (Lakshminarayanan et al., 2017). Combined with their relative simplicity, these properties make ensembling the method of choice for many applications, such as medical imaging and autonomous driving, where reliability is paramount.
Despite these advantages, the process of selecting post-hoc models that are both accurate and diverse remains a challenging combinatorial problem, especially as the pool of candidate models grows. Commonly used heuristics, particularly in the context of tabular data, such as greedy selection (Caruana et al., 2004) and various weighting schemes, attempt to optimize ensemble performance based on metrics evaluated on a held-out validation set or through cross-validation. However, these methods face significant limitations. Specifically, the selection of models to include in the ensemble and the determination of optimal ensembling strategies (e.g., stacking weights) are critical decisions that, if not carefully managed, can lead to overfitting on the validation data. Although neural networks are good candidates for generating ensembling weights, few studies rely on them as a post-hoc ensembling approach. We believe this happens primarily due to a lack of ensemble-related inductive biases that provide regularization.
In this work, we introduce a novel approach to post-hoc ensembling using neural networks. Our proposed Neural Ensembler dynamically generates the weights for each base model in the ensemble on a per-instance basis, a.k.a dynamical ensemble selection (Ko et al., 2008). To mitigate the risk of overfitting the validation set, we introduce a regularization technique inspired by the inductive biases inherent to the ensembling task. Specifically, we propose randomly dropping base models during training, inspired by previous work on DropOut in Deep Learning (Srivastava et al., 2014).
@@
    We propose a simple yet effective post-hoc ensembling method based on a neural network that dynamically ensembles base models.
@@
    To prevent the formation of low-diversity ensembles, we introduce a regularization technique that involves randomly dropping base model predictions during training. We demonstrate theoretically that this lower bounds the diversity of the generated ensemble. Additionally, we ablate its effect through various experiments.
@@
To promote reproducibility, we have made our code publicly available in the following anonymous repository  [^1] https://github.com/machinelearningnuremberg/NeuralEnsemblers/. We hope that our codebase, along with the diverse set of benchmarks used in our experiments, will serve as a valuable resource for the development and evaluation of future post-hoc ensembling methods.
@@ 2 Background and Motivation
Post-Hoc ensembling uses set of fitted base models $\{z_{1},...,z_{M}\}$ such that every model outputs $z_{m}(x):\mathbb{R}^{D}\rightarrow\mathbb{R}$. These outputs are combined by a stacking ensembler $f(z(x);\theta):=f(z_{1}(x),...,z_{M}(x);\theta):\mathbb{R}^{M}\rightarrow\mathbb{R}$, where $z(x)=[z_{1}(x),...,z_{M}(x)]$ is the concatination of the base models predictions. While the base models are estimated using a training set $\mathcal{D}_{\mathrm{Train}}$, the ensembler’s parameters $\theta$ are typically obtained by minimizing a loss function on a validation set $\mathcal{D}_{\mathrm{Val}}$ such that:
@@
In the general case, this objective function can be optimized using gradient-free optimization method such as evolutionary algorithms (Purucker & Beel, 2023b) or greedy search (Caruana et al., 2004) together with a linear combination $\theta\in\mathbb{R}^{M}$ of the model outputs:
@@
which is referred to as Bayesian Model Average, and uses $\theta_{m}=p(m)$. In the general case, the probabilistic ensembler $p(y|x)=p(y|z_{1}(x),...,z_{M}(x),\beta)$ is a stacker model parametrized by $\beta$.
@@ 2.1 Motivating Dynamic Ensembling
We motivate in this Section the need for dynamic ensembling by analyzing base models’ predictions in real data taken from our experimental metadatasets. In reality, Equation 3 does not specify the distribution $p(m)$. Generally, it is safe to assume that the performance associated with an ensembler $f(z_{m}(x),\theta)$ is optimal if we dynamically select the optimal aggregation $\theta_{m}(x)=p(m|m)$ on a per-data point basis, instead of a static $\theta$. To motivate this observation, we selected four datasets from different modalities: TabRepo (Tabular data (Salinas & Erickson, 2023)), QuickTune (Computer Vision (Arango et al., 2024)), FTC (NLP-Section D.1) and NasBench 201 (NAS for Computer Vision (Dong & Yang, 2020)). Then, we compute the per-sample error for 100 models in 100 samples. We report the results in Figure 1, indicating failed predictions with dark colors. We observe that models make different errors across samples, demonstrating the lack of optimality for static ensembling weights.
@@ 3 Neural Ensemblers
We use neural networks as ensemblers by training on the base model predictions $z(x)=[z_{1}(x);...;z_{M}(x)]$ as input. For regression, $z(x):\mathbb{R}^{D}\rightarrow\mathbb{R}^{M}$ outputs thebase models’ point predictions given by $x\in\mathbb{R}^{D}$, while for classification $z(x;c):\mathbb{R}^{D}\rightarrow[0,1]^{M}$ returns the probabilities predicted by the base models for class $c$. In our discussion we consider two functional modes for the ensemblers: as network outputting weights for model averaging or as a direct stacking model that outputs the prediction. In stacking mode for regression, we aggregate the base model point predictions using a neural network to estimate the final prediction $\hat{y}=\phi(z;\beta)$, where $\beta$ are the network parameters. In the model-averaging mode, the Neural Ensembler outputs the weights $\theta_{m}(z;\beta)$ to combine the model predictions as in Equation 4.
$\hat{y}=\sum_{m}\theta_{m}({z};\beta)\cdot z_{m}(x).$ (4)
@@
Regardless of the functional mode, the Neural Ensembler has a different output $\hat{y}$ for regression and classification. In regression, the output $\hat{y}$ is a point estimation of the mean for a normal distribution such that $p(y|x;\beta)=\mathcal{N}(\hat{y},\sigma)$. For classification, the input is the probabilistic prediction of the base models per class $z_{m}(x;c)=p(y=c|x,m)$, while the output is a categorical distribution $\hat{y}=p(y=c|z(x),\beta)$. We optimize $\beta$ by minimizing the negative log-likelihood over the validation dataset $\mathcal{D}_{\mathrm{Val}}$ as:
$\min_{\beta}\mathcal{L}(\beta;D_{\mathrm{Val}})=\min_{\beta}\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}-\mathrm{log}\ p(y|x;\beta).$ (5)
3.1 Architecture
@@
We discuss the architectural implementation of the Neural Ensembler for the classification case, which we show in Figure 2. For the stacking mode, we use an MLP that outputs the logit $\phi(z(x;c);\beta):\mathbb{R}^{M}\rightarrow[0,1]$ of each class $c$. Specifically, the network receives as inputs the base models’ predictions $z(x,c)$ for the class $c$ and outputs the corresponding predicted logit for this class, i.e. $\phi(z;c)$. The model predictions per class are fed independently, enabling sharing the network parameters $\beta$. Subsequently, we compute the probability $p(y=c|x)=\frac{\exp^{\phi(z;c)}}{\sum_{c^{\prime}}\exp^{\phi(z;c^{\prime})}}$, with $\mathrm{MLP}(z(x;c);\beta)$. In regression, the final prediction is the output $\phi(z(x);\beta)$.
For model averaging mode, we use a novel architecture based on a Deep Set (Zaheer et al., 2017) embedding of the base models predictions. We compute the dynamic weights $\theta_{m}(z;\beta)=\frac{\exp f_{m}(z;\beta)}{\sum_{m^{\prime}}f_{m^{\prime}}(z;\beta)}$, where the unnormalized weight per model $f_{m}(z;\beta)$ is determined via two MLPs. The first one $\text{MLP}_{1}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ embeds the predictions per class $z(x,c^{\prime})$ into a latent dimension of size $H$, whereas the second network $\text{MLP}_{2}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ aggregates the embeddings and outputs the unnormalized weights, as shown in Equation 6. Notice that the Neural Ensemblers’ input dimension and number of parameters are independent of the number of classes, due to our proposed parameter-sharing mechanism.
$f(z;m,\beta)=\mathrm{MLP}_{2}\left(\sum_{c^{\prime}}\mathrm{MLP}_{1}\left(z(x;c^{\prime});\beta_{1}\right);\beta_{2}\right).$ (6)
3.2 The Risk of Diversity Collapse
In this section, we elaborate on one risk that might arise during learning ensembles, as a base model might be highly correlated with the target in the validation set, when building ensembles. We dub this model a Preferred Base Model.
Definition 1 (Preferred Base Model).
Proposition 1.
If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{p}\}$, given the conditions of Definition 1.
Sketch of Proof.
Given that the base models predictions are uncorrelated, their correlations $\rho_{z_{m},z_{p}}=0,\forall m\neq p$, then it holds that $-\sqrt{1-\rho_{z_{p},y}^{2}}\leq\rho_{z_{m},y}\leq\sqrt{1-\rho_{z_{p},y}^{2}},\forall z_{m}\in\mathcal{F}/\{z_{p}\}$. We provide more details in the Appendix A. ∎
@@ 3.2 The Risk of Diversity Collapse
On the other hand, an important aspect when building ensembles is guaranteeing diversity among the models (Wood et al., 2023; Jeffares et al., 2024). This has motivated some approaches to explicitly account for diversity when searching the ensemble configuration  (Shen et al., 2022; Purucker et al., 2023). A common way to measure diversity is the ambiguity (Krogh & Vedelsby, 1994), which can be derived after decomposing the loss function (Jeffares et al., 2024). Unfortunately, even simple ensembles risk obtaining low diversity although their base models are uncorrelated. As we will shortly show, this happens especially when there is a preferred base model.
Proposition 2 (Diversity Collapse).
If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$, i.e. $\lim_{\rho_{z_{p},y}\to 1}\alpha=0$.
We give a detailed proof in the Appendix A. However, we can build the intuition after noticing that given Proposition 1, $\lim_{\rho_{z_{p},y}\to 1}\bar{z}=z_{p}$.
@@ 3.3 Base Models’ DropOut
6      Sample masking vector $r\in\mathbb{R}^{M},r_{i}\sim\mathrm{Ber}(\gamma)$;
@@
8      Mask base models predictions $z_{\mathrm{drop}}(x)=r\odot z(x)$ ;
@@
11             Compute predictions $\hat{y}=\phi\left(\frac{1}{\gamma}z_{\mathrm{drop}};\beta\right)$ ;
@@
13
14            Compute weights $\theta\left(\frac{1}{\gamma}z_{\mathrm{drop}};m,\beta,r\right)$ using Equation 7 ;
15
16            Compute predictions $\hat{y}$ using Equation 4 ;
17
@@ 16            Compute predictions $\hat{y}$ using Equation 4 ;
19      Update neural ensembles parameters $\beta$ using $\nabla\mathcal{L}(\beta;\mathcal{D}_{\mathrm{Val}})$
@@
Using Neural Ensemblers tackles the need for dynamical ensembling. Moreover, it gives additional expressivity associated with neural networks. However, there is also a risk of overfitting and diversity collapse. As we showed in Section 3.2, this happens when there is a preferred model on which the ensembler mainly relies while neglecting other base model predictions. Although it might effectively decrease the validation loss (Equation 5), it does not necessarily generalize to test samples. Inspired by previous work (Srivastava et al., 2014), we propose to drop some base models during training forward passes. Intuitively, this forces the ensembler to rely on different base models to perform the predictions, instead of merely using the preferred base model(s).
Formally, we mask the inputs such as $r_{m}\cdot z_{m}(x)$, where $r_{m}\sim\mathrm{Ber}(\gamma)$, where $\mathrm{Ber}(\gamma)$ is the Bernoulli distribution with parameter $\gamma$ with represents the probability of keeping the base model, while $\delta=1-\gamma$ represents the DropOut rate. We also mask the weights when using model averaging:
$\theta(z;m,\beta,r)=\frac{r_{m}\cdot\exp{f_{m}}({z};\beta)}{\sum_{m^{\prime}}r_{m^{\prime}}\cdot\exp{f_{m^{\prime}}}({z};\beta)}.$ (7)
As DropOut changes the scale of the inputs, the weight scaling rule should be applied during inference by multiplying the dropped variables by the retention probability $\gamma$. Alternatively, we can scale the variables during training by $\frac{1}{\gamma}$. In Algorithm 1, we detail how to train the Neural Ensemblers by dropping base model predictions. It has two modes, acting as a direct stacker or as a model averaging ensembler. We demonstrate that base models’ DropOut avoids diversity collapse by upper bounding the diversity as stated in Proposition 6, even for the simplest ensembling case.
Proposition 3 (Avoiding Diversity Collapse).
Sketch of Proof. We want to compute $\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$. By using $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$, and assuming, without loss of generality, that the predictions are standardized, we obtain $\mathbb{V}(r\cdot z_{m})=\gamma$. This lead as to $\lim_{\rho_{z_{p},y}\to 1}\alpha=1-\gamma$, after following a procedure similar to Proposition 2. We provide the complete proof in Appendix A.

### v2

# Regularized Neural Ensemblers

@@ Abstract
Ensemble methods are known for enhancing the accuracy and robustness of machine learning models by combining multiple base learners. However, standard approaches like greedy or random ensembling often fall short, as they assume a constant weight across samples for the ensemble members. This can limit expressiveness and hinder performance when aggregating the ensemble predictions. In this study, we explore employing regularized neural networks as ensemble methods, emphasizing the significance of dynamic ensembling to leverage diverse model predictions adaptively. Motivated by the risk of learning low-diversity ensembles, we propose regularizing the ensembling model by randomly dropping base model predictions during the training. We demonstrate this approach provides lower bounds for the diversity within the ensemble, reducing overfitting and improving generalization capabilities. Our experiments showcase that the regularized neural ensemblers yield competitive results compared to strong baselines across several modalities such as computer vision, natural language processing, and tabular data.
@@ 1  Introduction
Ensembling machine learning models is a well-established practice among practitioners and researchers, primarily due to its enhanced generalization performance over single-model predictions (Ganaie et al.,, 2022; Erickson et al.,, 2020; Feurer et al.,, 2015; Wang et al.,, 2020). Ensembles are favored for their superior accuracy and ability to provide calibrated uncertainty estimates and increased robustness against covariate shifts (Lakshminarayanan et al.,, 2017). Combined with their relative simplicity, these properties make ensembling the method of choice for many applications, such as medical imaging and autonomous driving, where reliability is paramount. A popular set-up consists of ensembling from a pool, after training them separately, a.k.a. post-hoc ensembling (Purucker and Beel, 2023b, ). This allows users to use models trained during hyperparameter optimization.
Despite these advantages, selecting post-hoc models that are accurate and diverse remains a challenging combinatorial problem, especially as the pool of candidate models grows. Commonly used heuristics, particularly in the context of tabular data, such as greedy selection (Caruana et al.,, 2004) and various weighting schemes, attempt to optimize ensemble performance based on metrics evaluated on a held-out validation set or through cross-validation. However, these methods face significant limitations. Specifically, choosing models to include in the ensemble and determining optimal ensembling strategies (e.g., stacking weights) are critical decisions that, if not carefully managed, can lead to overfitting on the validation data. Although neural networks are good candidates for generating ensembling weights, few studies rely on them as a post-hoc ensembling approach. We believe this happens primarily due to a lack of ensemble-related inductive biases that provide regularization.
In this work, we introduce a novel approach to post-hoc ensembling using neural networks. Our proposed Neural Ensembler dynamically generates the weights for each base model in the ensemble on a per-instance basis, a.k.a dynamical ensemble selection (Ko et al.,, 2008; Cavalin et al.,, 2013). To mitigate the risk of overfitting the validation set, we introduce a regularization technique inspired by the inductive biases inherent to the ensembling task. Specifically, we propose randomly dropping base models during training, inspired by previous work on DropOut in Deep Learning (Srivastava et al.,, 2014). Furthermore, our method is modality-agnostic, as it only relies on the base model predictions. For instance, for classification, we use the class predictions of the base models as input.
@@
    We propose a simple yet effective post-hoc ensembling method based on a regularized neural network that dynamically ensembles base models and is modality-agnostic.
@@
    To prevent the formation of low-diversity ensembles, the regularization technique randomly drops base model predictions during training. We demonstrate theoretically that this lower bounds the diversity of the generated ensemble, and validate its effect empirically.
@@
To promote reproducibility, we have made our code publicly available in the following anonymous repository [^1] https://github.com/machinelearningnuremberg/RegularizedNeuralEnsemblers. We hope that our codebase, along with the diverse set of benchmarks used in our experiments, will serve as a valuable resource for the development and evaluation of future post-hoc ensembling methods.
@@ 2  Background and Motivation
Post-hoc ensembling uses set of fitted base models $\{z_{1},...,z_{M}\}$ such that every model outputs predictions $z_{m}(x):\mathbb{R}^{D}\rightarrow\mathbb{R}$. These outputs are combined by a stacking ensembler $f(z(x);\theta):=f(z_{1}(x),...,z_{M}(x);\theta):\mathbb{R}^{M}\rightarrow\mathbb{R}$, where $z(x)=[z_{1}(x),...,z_{M}(x)]$ is the concatenation of the base models predictions. While the base models are fitted using a training set $\mathcal{D}_{\mathrm{Train}}$, the ensembler’s parameters $\theta$ are typically obtained by minimizing a loss function on a validation set $\mathcal{D}_{\mathrm{Val}}$ such that:
@@
In the general case, this objective function can be optimized using gradient-free optimization methods such as evolutionary algorithms (Purucker and Beel, 2023b, ) or greedy search (Caruana et al.,, 2004). Commonly, $f$ is a linear combination $\theta\in\mathbb{R}^{M}$ of the model outputs:
@@
which is referred to as Bayesian Model Average (Raftery et al.,, 1997), and uses $\theta_{m}=p(m)$. In the general case, the probabilistic ensembler $p(y|x)=p(y|z_{1}(x),...,z_{M}(x),\beta)$ is a stacking model parametrized by $\beta$.
@@ 2.1  Motivating Dynamic Ensembling
We motivate in this Section the need for dynamic ensembling by analyzing base models’ predictions in real data taken from our experimental meta-datasets. Generally, the distribution $p(m)$ can take different forms. Dynamic ensembling assumes that the performance associated with an ensembler $f(z_{m}(x),\theta)$ is optimal if we select the optimal aggregation $\theta_{m}(x)=p(x|m)$ on a per-data point basis, instead of a static $\theta$. To motivate this observation, we selected four datasets from different modalities: TabRepo (Tabular data (Salinas and Erickson,, 2023)), QuickTune (Computer Vision (Arango et al., 2024a, )), FTC (Arango et al., 2024b, ) and NasBench 201 (NAS for Computer Vision (Dong and Yang,, 2020)). Then, we compute the per-sample error for 100 models in 100 samples. We report the results in Figure 1, indicating failed predictions with dark colors. We observe that models make different errors across samples, indicating a lack of optimality in static ensembling weights.
@@ 3  Neural Ensemblers
[…]
$\sum_{m}\theta_{m}({z};\beta)\cdot z_{m}$ (4)
@@
Regardless of the functional mode, the Neural Ensembler has a different output $\hat{y}$ for regression and classification. In regression, the output $\hat{y}$ is a point estimation of the mean for a normal distribution such that $p(y|x;\beta)=\mathcal{N}(\hat{y},\sigma)$. For classification, the input is the probabilistic prediction of the base models per class $z^{(c)}_{m}=p(y=c|x,m)$, while the output is a categorical distribution $\hat{y}=p(y=c|z,\beta)$. We optimize $\beta$ by minimizing the negative log-likelihood over the validation dataset $\mathcal{D}_{\mathrm{Val}}$ as:
$\min_{\beta}\mathcal{L}(\beta;D_{\mathrm{Val}})=\min_{\beta}\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}-\mathrm{log}\ p(y|x;\beta)$ (5)
3.1  An Architecture with Parameter Sharing
@@
We discuss the architectural implementation of the Neural Ensembler for the classification case, which we show in Figure 2. For the stacking mode, we use an MLP that receives as input the base models’ predictions $z^{(c)}$ for the class $c$ and outputs the corresponding predicted logit for this class, i.e. $\phi(z,c;\beta)$. The model predictions per class are fed independently into the same network, enabling the sharing of the network parameters $\beta$ as shown in Figure 6. Subsequently, we compute the probability $p(y=c|x)=\frac{\exp^{\phi(z,c;\beta)}}{\sum_{c^{\prime}}\exp^{\phi(z,c^{\prime};\beta)}}$, with $\phi(x,c;\beta)=\mathrm{MLP}(z^{(c)};\beta)$. In regression, the final prediction is the output $\phi(z;\beta)$.
[…]
$f_{m}(z;\beta)=\mathrm{MLP}_{2}\left(\sum_{c^{\prime}}\mathrm{MLP}_{1}\left(z^{(c^{\prime})};\beta_{1}\right);\beta_{2}\right)$ (6)
3.2  Ensemble Diversity
@@
An important aspect when building ensembles is guaranteeing diversity among the models (Wood et al.,, 2023; Jeffares et al.,, 2024). This has motivated some approaches to explicitly account for diversity when searching the ensemble configuration  (Shen et al.,, 2022; Purucker et al.,, 2023). A common way to measure diversity is the ambiguity (Krogh and Vedelsby,, 1994), which can be derived after decomposing the loss function (Jeffares et al.,, 2024). Measuring diversity is essential as it helps to avoid overfitting in the ensemble.
Definition 3.1 (Ensemble Diversity via Ambiguity Measure).
[…]
@@ 3.3  Base Models’ DropOut
[…]
@@
8      Mask base models predictions $z_{\mathrm{drop}}=r\odot z$ ;
@@
11             Compute predictions $\phi\left(\frac{1}{\gamma}z_{\mathrm{drop}}\right)$ ;
12
@@ 11             Compute predictions $\phi\left(\frac{1}{\gamma}z_{\mathrm{drop}}\right)$ ;
14
15            Compute weights $\theta_{m}\left(\frac{1}{\gamma}z_{\mathrm{drop}}\right)$ using Equation 7 ;
16
17            Compute predictions using Equation 4 ;
18
@@ 17            Compute predictions using Equation 4 ;
20      Update parameters $\beta$ using $\nabla\mathcal{L}(\beta;\mathcal{D}_{\mathrm{Val}})$
@@
Using Neural Ensemblers tackles the need for dynamical ensembling. Moreover, it gives additional expressivity associated with neural networks. However, there is also a risk of overfitting caused by a low diversity. Although the error might effectively decrease the validation loss (Equation 5), it does not necessarily generalize to test samples. Inspired by previous work (Srivastava et al.,, 2014), we propose to drop some base models during training forward passes. Intuitively, this forces the ensembler to rely on different base models to perform the predictions, instead of merely using the preferred base model(s).
Formally, we mask the inputs such as $r_{m}\cdot z_{m}$, where $r_{m}\sim\mathrm{Ber}(\gamma)$, where $\mathrm{Ber}(\gamma)$ is the Bernoulli distribution with parameter $\gamma$ with represents the probability of keeping the base model, while $\delta=1-\gamma$ represents the DropOut rate. We also mask the weights when using model averaging:
$\theta_{m}(z;\beta,r)=\frac{r_{m}\cdot\exp{f_{m}}({z};\beta)}{\sum_{m^{\prime}}r_{m^{\prime}}\cdot\exp{f_{m^{\prime}}}({z};\beta)}.$ (7)
As DropOut changes the scale of the inputs, we should apply the weight scaling rule during inference by multiplying the dropped variables by the retention probability $\gamma$. Alternatively, we can scale the variables during training by $\frac{1}{\gamma}$. In Algorithm 1, we detail how to train the Neural Ensemblers by dropping base model predictions. It has two modes, acting as a direct stacker or as a model averaging ensembler. We demonstrate that base models’ DropOut lower bounds the diversity for a simple ensembling case in Proposition A.1.
Definition 3.2 (Preferred Base Model).
Proposition 3.3 (Diversity Lower Bound).
As the correlation of the preferred model increases $\rho_{p_{m},y}\rightarrow 1$, the diversity decreases $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
[…]